OpenStudy (anonymous):

@phi The moon forms a right triangle with the Earth and the Sun during one of its phases, as shown below: (A triangle with sides Earth, Sun, and Moon. Side Earth-Sun has distance y. Angle x is on the same side) A scientist measures the angle x and the distance y between the Earth and the Sun. Using complete sentences, explain how the scientist can use only these two measurements to calculate the distance between the moon and the Sun.

OpenStudy (phi):

can you post or draw the picture?

OpenStudy (phi):

if we call the points M,E and S where is the right angle? which side is the hypotenuse ?

OpenStudy (anonymous):

Sorry, computer slow.

OpenStudy (anonymous):

Sorry, computer slow.

OpenStudy (phi):

this looks like trig. you know angle x and side y you want to know the side opposite angle x do you know SOH CAH TOA ?

OpenStudy (anonymous):

Sorry, computer slow.

OpenStudy (anonymous):

yes, its just that when the angle is on is on the same side as Y it is throwing me off on which function to use.

OpenStudy (phi):

do you know how to "name" the sides? which side is the hypotenuse, and standing at angle x, which side is 1) adjacent 2) opposite ?

OpenStudy (anonymous):

ME is the opposite ES is the adjacent MS is the hypotenuse

OpenStudy (phi):

they say they "measured side y" which is ES = adj you want to find ME= opp so I would use the trig function that uses opposite and adjacent

OpenStudy (anonymous):

Actually it wants me to find MS if I am not mistaken, says to use given information (x and y) to find the distance between the moon and the sun.

OpenStudy (phi):

good point. MS is hypotenuse and you know ES = adj so use the trig function that uses hyp and adj

OpenStudy (anonymous):

Cosine or Secant.

OpenStudy (phi):

nobody uses secant these days, except for specialized reasons

OpenStudy (anonymous):

Okay, so I would start of with Cos = Adj/Hyp Cos x = y/x

OpenStudy (phi):

x is used for the angle, so I would use a different letter for the unknown side. how about d (for distance) ?

OpenStudy (anonymous):

so y/d?

OpenStudy (phi):

as long as you know d= MS cos x = y/d solve for d

OpenStudy (anonymous):

so, Cos = Adj/Hyp Cos x = y/d x = d

OpenStudy (phi):

Cos x = y/d to solve for d, you can multiply both sides by d d cos x = y*d/d the right side simplifies to y*1 or just y d cos x = y now divide both sides by cosx d = y/cos x

OpenStudy (anonymous):

So how would this look as a whole? Just making sure I getting it.

OpenStudy (phi):

try writing up what you think the answer is

OpenStudy (anonymous):

Cos = Adj/Hyp Cos x = y/d (or y*d/d, having trouble with this part) Cos x = y d = y/cos x

OpenStudy (anonymous):

or Cos = Adj/Hyp Cos x = y/d x (x) = d x= d

OpenStudy (phi):

start by saying The given right triangle with angle x, has measured side y (adjacent side), and unknown side d (side MS, or hypotenuse) thus we should use CAH, Cos x= Adj/Hyp Cos x = y/d the next step is algebra \[ \cos x = \frac{y}{d} \] if two sides are equal, and we multiply both sides by the same thing, they stay equal in other words, if we multiply both sides by "d", we still have an equation \[ d \cdot \cos x = d \cdot \frac{y}{d} \] however \( d \cdot \frac{y}{d} \) is the same thing as \( \frac{d \cdot y}{y} \) and that is the same thing as \( y \cdot \frac{d}{d} \) and that is the same as \( y \cdot 1\) or just y

OpenStudy (phi):

so \[ d \cos x = y \] now we want d by itself. Divide both sides by cos x \[ d \frac{\cos x}{\cos x} = \frac{y}{\cos x} \] and again, something divided by itself is 1, so \[ d= \frac{y}{\cos x} \]

OpenStudy (anonymous):

thanks your help! can you check my answers to other problem (just three)